40 grams of liquid A and 70 grams of liquid B are mixed together. If both are ideal solutions, calculate the vapor composition. p0A = 200 mm, p0B = 700 mm, МA = 56 gr mol-1, MB = 90 gr mol-1.

Grade:12th pass

2 Answers

Soumendu Majumdar
5 years ago
Dear Vira,Accordin
Soumendu Majumdar
5 years ago
Dear Vira,
According to Raoult’s Law,
P= P0\chi & PB = P0B \chiB      where \chiA & \chiB are mole fractions of liquids A & B respectively.
Number of moles of A =40\div56 mol
Number of moles of B =70\div90 mol
Mole fraction of A =(40/56)\div[(40/56)+(70/90)]=0.47872340425
Mole fraction of B =(70/90)\div[(40/56)+(70/90)]=0.52127659574
You may round off the values such that  \chiA + \chi= 1 remains true.
Now using Raoult’s Law, If the total vapor pressure of the liquids be P then;
P = PA + P B
=[(200\times 0.47872340425) +(700\times 0.52127659574)] mm
=460.638297868 mm
You may approximate it to 460.6 mm or any other relevant value as per your need.
Hope this helps!
regards,
Soumendu